OpenStudy (anonymous):

\[\frac{ \sqrt{90} }{ \sqrt{10} }\]

zepdrix (zepdrix):

Hey there Ash :) Need some help simplifying/rationalizing?

OpenStudy (anonymous):

I'm not sure all it says is evaluate the expression I'm not sure how to do it though

zepdrix (zepdrix):

We can use this approach to rewrite our numerator:\[\Large\rm \sqrt{90}=\sqrt{9\cdot10}=\sqrt9\cdot\sqrt{10}\] \[\Large\rm \frac{\color{orangered}{\sqrt{90}}}{\sqrt{10}}=\frac{\color{orangered}{\sqrt{9}\cdot\sqrt{10}}}{\sqrt{10}}\]Hmm do you see how that helps us?

zepdrix (zepdrix):

Can we "cancel" some stuff out maybe?

OpenStudy (anonymous):

What do you mean?